# A Data-Centric Approach To Generate Faithful and High Quality Patient Summaries with Large Language Models

### v1

@@ 4.1 Generation of Patient Summaries
4.2 Protocol for Labeling Hallucinations
We developed a protocol for labeling hallucinations in patient summaries based on (Thomson and Reiter, 2020, 2021). [^7] See word document on Github: https://github.com/ehudreiter/accuracySharedTask/blob/main/example_exercise/Example_Annotation_Exercise.docx We created similar instructions with adjustments for our use case. First, we selected labels that were more relevant for patient summaries. Since most hallucinations were unsupported facts, we further distinguished them into nine subcategories. Contradicted facts and incorrect facts were kept as general error labels (see legend in Figure 2). Second, Thomson and Reiter (2020) considered facts from all sources as given, i.e., annotators could also use information from the internet to check facts in the summary. In contrast to that, we treated the context (BHC) as the only ground truth about the patient. We chose this approach to reduce the labeling burden as annotators could not be expected to review all notes and structured information of a patient. However, since patient summaries do not only contain patient specific information, we did allow general medical knowledge and advice even if not explicitly provided in the context (e.g., “Please take your medications as prescribed”). Our labeling protocol is attached at the end of the Appendix and can be downloaded from our Github repository.
@@ 4.3 Hallucinations in MIMIC and Generated Patient Summaries
We used our developed protocol to annotate hallucinations in the original MIMIC and generated patient summaries. It is important to note, that “hallucinations” in the MIMIC summaries differ from model-generated hallucinations. Unsupported or contradicted facts in the summaries written by doctors can be due to several reasons. Doctors may include information in the summary that was never documented, that was documented outside the considered context (in our case, only the BHC), or that was altered just prior to discharge. Therefore, the presence of hallucinations in patient summaries is common in healthcare practice and should usually not be regarded as real errors. In this work, we considered the MIMIC data from a machine learning standpoint and analyzed the effect of hallucinations in the training data on LLMs.
@@
MIMIC / Number of Annotated Hallucinations per Type
@@
Table 3: The first row shows the labeling results for 100 MIMIC summaries for which we found 286 hallucinations. The remaining rows show the category breakdown for different types of hallucinations annotated on 20 summaries. It corresponds to the data in the Hallucinationsnormal-$\downarrow$ column in Table 2.
First, we labeled 100 MIMIC context-summary pairs (see Figure 2) to characterize the quantity and types of hallucinations in the data. Based on this, we created the cleaned and the cleaned & improved datasets, each containing 100 examples (see Section 3.2). We also labeled hallucinations in 100 generated summaries, which were generated using five different models for 20 examples (see Section 4.4). This was done to determine a potential hallucinations reduction. To prevent model identification, we permuted the five generated summaries for each example.
The labeling was carried out by two German medical students in their sixth year. They had completed their second state examination (equivalent to USMLE Step 2) and were working in the hospital. We utilized MedTator for annotation (He et al., 2022). For training, both annotators labeled two times five examples that were discussed and agreed upon. For the final labeling, the annotators worked independently and reached a consensus through discussion. We determined manual and automatic statistics for agreement (see Appendix B). Clarifications regarding the labeling procedure that arose during training or the final annotation were added to the protocol.
4.4 Data-Centric Hallucination Reduction
We evaluated whether manually removing hallucinations from the training data can reduce hallucinations generated by LLMs. To this end, we tested Llama 70B and GPT-4 5-shot trained or prompted on the original dataset and the hallucination-free cleaned dataset (see Section 3.2). For Llama 70B, we fine-tuned on all 100 examples, while for GPT-4 5-shot, we randomly selected 5 examples. We also included GPT-4 0-shot in this experiment, which did not require any training data. Figure 9 contains example generations from all models. Due to limited manual annotation capabilities, no further models were included in this analysis. Using these five models, we generated summaries for 20 examples. We labeled hallucinations in these 100 generated summaries as described in Section 4.3. We also determined the amount of missing key facts and medical jargon. For this, we employed the same annotation procedures used for the qualitative evaluation (see Appendix F).
Figure 3: Generated summaries from Llama 70B trained on 100 original and 100 cleaned examples for synthetic context given in Figure 2 with manually labeled hallucinated spans highlighted (see Figure 2 for color map). These are two of the five models included for data-centric hallucination reduction for which we annotated hallucinations in generated summaries.
@@ 4.5 Quantitative Evaluation
We performed a quantitative performance evaluation for all models to compare them to existing work and analyze the influence of the number of training examples. We used data from MIMIC-IV-Note-DI (MIMIC) and used all examples (LED), 100 examples (Llama), or 5 examples (GPT-4 5-shot) for training. During parameter tuning (see Section C), we used 100 examples each for validation and testing.
To evaluate lexical overlap, we used the F1 score of ROUGE (Lin, 2004). For similarity based on contextual embeddings, we reported BERTScore (Zhang et al., 2020a) for the default roberta-large (BERTScore) and microsoft/deberta-large-mnli (DeBERT) as recommended by the authors. [^8] See readme: https://github.com/Tiiiger/bert_score We did not utilize medical embeddings since the summaries should be written in layperson language. Lastly, we determined the SARI score for text simplification (Xu et al., 2016) and the number of generated words.
@@ 4.6 Qualitative Evaluation
We evaluated the quality of patient summaries generated by LLMs that were trained or prompted using the highest quality data (cleaned & improved). We used all 100 examples to fine-tune LLama 70B and sampled 5 random examples for GPT-4 5-shot. We also included the original MIMIC summaries, LED-large trained on all MIMIC data, and GPT-4 0-shot. Figure 10 shows example generations from all models.
@@
We evaluated the generated summaries for Relevance, Consistency, Fluency, and Coherence using a 1-5 Likert scale (Fabbri et al., 2021) following previous work on medical summarization (Adams et al., 2022). We added the additional dimension Simplification. To ensure more reliable results, we defined subtasks to be performed before rating relevance (subtask: labeling key facts in context and summary), consistency (labeling hallucinations using our protocol), and simplification (labeling medical jargon). Also, we provided clear definitions for each Likert rating. Additional details on the subtasks for labeling key facts and medical jargon and annotator instructions are given in Appendix F and Figures 7 and 8.
The annotators who labeled hallucinations also conducted the qualitative evaluation. To minimize their workload, we used the same 20 examples. We again permuted the generated summaries to prevent model identification. The annotators completed two training examples, and no agreement was sought as this was considered a subjective task. We combined both ratings for our analysis.
Model (training data) R-1normal-$\uparrow$ R-2normal-$\uparrow$ R-3normal-$\uparrow$ R-4normal-$\uparrow$ R-Lnormal-$\uparrow$ BERTnormal-$\uparrow$ DeBERTnormal-$\uparrow$ SARInormal-$\uparrow$ Words
LED-large (all MIMIC ex.) $43.82$ $17.33$ $8.85$ $4.92$ $29.89$ $88.11$ $64.12$ $46.71$ $76.99$
Llama 2 7B (100 MIMIC ex.) $38.36$ $12.66$ $5.13$ $2.24$ $24.73$ $85.68$ $60.23$ $44.12$ $73.13$
Llama 2 70B (100 MIMIC ex.) $40.58$ $14.31$ $6.09$ $2.74$ $26.19$ $86.30$ $61.89$ $45.16$ $76.90$
GPT-4 5-shot (5 MIMIC ex.) $38.80$ $10.78$ $3.55$ $1.12$ $21.98$ $86.67$ $61.30$ $42.88$ $131.86$
@@
Table 4: Quantitative evaluation for patient summary generation on the MIMIC-IV-Note-DI dataset. We used the ROUGE F1-score for n-grams (R-n) and the longest common subsequence (R-L), BERTScore using roberta-large (BERT) and deberta-large-mnli (DeBERT), the SARI score, and the number of generated words. Additional performance results can be found in Table 12.
4.7 Automatic Hallucination Detection
Based on annotated hallucinations in 100 MIMIC and 100 generated summaries, we wanted to evaluate automatic hallucination detection. This can be considered as span classification: for a given context $C$ and a summary $S$, we aim to identify all hallucination spans in a summary $\mathcal{H}_{S}=\{H_{1},...,H_{n}\}$ with $H_{i}=(s,e,c)$ consisting of a start token, end token, and a class (see Figure 2). We also explored class-agnostic hallucination detection without class $c$.
We tested an approach based on medical entities for class-agnostic hallucination detection (Cai et al., 2022; Adams et al., 2022). We used MedCat (Kraljevic et al., 2021) for entity recognition and linking to UMLS in the context and summary. We used the semantic UMLS types from Adams et al. (2021). We considered all entities that appeared in the summary but not in the context as hallucinations. To enhance this approach, we incorporated SapBERT embeddings (Liu et al., 2021) to determine entity equivalence (Adams et al., 2023). The equivalence threshold for cosine similarity was determined based on ten examples labeled during annotator training. Furthermore, we tested GPT-4 for class-agnostic and class-aware hallucination detection. We designed a suitable prompt based on the aforementioned ten examples and utilized varying numbers of in-context examples and chain-of-thought (COT) prompting, detailed in Appendix E. Evaluation was performed using partial matching with the nervaluate package (Segura-Bedmar et al., 2013).
@@ 5.1 Hallucinations in MIMIC and Generated Patient Summaries
We labeled 100 MIMIC summaries and found 286 hallucinations (see the first row in Table 3). The most prevalent label was the generic word unsupported (n=76) followed by condition unsupported (n=52), and time unsupported (n=35). Hence, there is a significant amount of unsupported data in MIMIC when using the BHC as context. The annotation statistics show that annotator 2 labeled more than annotator 1 (2.82 versus 2.39) and they agreed on 1.55 of 2.86 annotations per summary (see Table 9).
For the 100 generated patient summaries, we found a total of 114 hallucinations and word unsupported was the most common label type (see Table 3). Annotator 1 labeled more hallucinations than annotator 2 (1.23 versus 1.18) and they agreed on 0.67 of 1.14 annotations per summary (see Table 9). The annotation of MIMIC summaries took 30 hours for each annotator and 6 hours for the agreement. For the generated summaries, annotation took less time, with 20 hours for labeling and 4 hours for the agreement, since there were only 20 different patient contexts.
@@ 5.2 Data-Centric Hallucination Reduction
The results for GPT-4 5-shot with original were significantly better than for Llama 70B trained on original with 0.70 versus 2.60 hallucinations and 0.93 versus 3.77 missed key facts per summary (see Table 2). Data-centric hallucination-reduction (training on cleaned), showed a substantial effect for Llama 70B, reducing hallucinations from 2.60 to 1.55 per summary while maintaining the same number of key facts. Figure 3 displays generations of both models with hallucination annotations. Llama 70B trained on original introduced unsupported adjectives “mild” and “other”, as well as an unsupported Metformin dosage. Both models incorrectly recommended continuing antibiotics at home. Apart from “mild”, all hallucination can be considered reasonable advice and are commonly found in patient summaries. For GPT-4 5-shot, the difference of 0.70 (original) to 0.40 (cleaned) hallucinations is only marginal, while retaining key facts. The use of medical jargon and the length of the generations did not change substantially. GPT-4 0-shot exhibited a few hallucinations (0.45) and the lowest number of missing key facts (0.82) and medical jargon (0.70), although the generations were longer than for 5-shot. Table 3 shows the types of hallucination for each model. While Llama 70B occasionally generated unsupported conditions and locations, this was rare for GPT-4. The hallucinations produced by GPT-4 were mostly labeled as unsupported word.
Figure 4: Qualitative evaluation for Llama 70B and GPT-4 5-shot trained and prompted on cleaned & improved data. We compared them to the MIMIC summaries, LED-large trained on all MIMIC data, and GPT-4 0-shot. 20 summaries were evaluated for each model by two medical experts.
@@ 5.3 Quantitative Evaluation
Table 4 contains the performance evaluation on the MIMIC-IV-Note-DI dataset with short context. Overall,the LED model performed best followed by Llama 2 and GPT-4. Only for BERT and DeBERT, Llama 2 and GPT-4 perform similarly. This observation aligns with the difference in training data. Notably, GPT-4 achieves considerable performance even without any training examples (GPT-4 0-shot).
@@ 5.4 Qualitative Evaluation
The results are visualized in Figure 4 and the mean ratings are given in Table 5. It should be noted that the consistency results for MIMIC summaries should be interpreted with caution because these summaries were artificially considered only with the restricted context of the BHC. The original MIMIC examples, LED-large trained on all data, and Llama 70B trained on 100 cleaned & improved examples showed similar results. On average, the relevance was rated at approximately two, indicating that three key points were missing on average. The simplification, fluency, and coherence were rated around four for all three approaches. Llama 70B performed better for consistency with a mean rating of 3.9 compared to 2.8 and 3.0 for the original and LED-large summaries, respectively. GPT-4 performed significantly better across all five dimensions. The difference in consistency compared to Llama 70B is supported by the lower hallucination rate of GPT-4 as observed earlier (see Table 2). Also, GPT-4 achieved high relevance ($\geq$ 4.0) with an average of one missing key fact per summary and average scores over 4.5 for simplification, fluency, and coherence. There was only a slight difference between GPT-4 5-shot and GPT-4 0-shot. GPT-4 5-shot performed better for consistency, i.e., including all key facts, while GPT-4 0-shot performed better for relevance and simplification.
Model (training data) Rel Con Sim Flu Coh
Original MIMIC $2.0$ $2.8$ $4.1$ $4.1$ $3.9$
LED-large (all MIMIC ex.) $2.1$ $3.0$ $3.7$ $3.5$ $3.4$
Llama 70B (100 cl.+imp. ex.) $1.9$ $3.9$ $3.9$ $3.8$ $3.6$
GPT-4 5-shot (5 cl.+imp. ex.) $4.0$ $4.7$ $4.5$ $4.9$ $5.0$
GPT-4 0-shot (none) $4.3$ $4.4$ $4.6$ $4.9$ $4.9$
Table 5: Mean qualitative ratings for original summaries, LED-large, Llama 70B and GPT-4 5-shot using cleaned & improved data (cl.+imp.), and GPT-4 0-shot [Relevance (Rel), Consistency (Con), Simplification (Sim), Fluency (Flu), Coherence (Coh)].
Model MIMIC Generated
@@
Table 6: Performance for hallucination detection on 100 examples from MIMIC-IV-Note-DI and 100 generated summaries. Ignored hallucination labels and used partial matching.

### v2

@@ 4.3 Hallucinations in MIMIC and Generated Patient Summaries
We used our developed protocol to annotate hallucinations in 100 doctor-written MIMIC summaries (Hallucinations-MIMIC-DI) and 100 LLM-generated summaries (Hallucinations-Generated-DI). It is important to note that “hallucinations” in doctor-written summaries are common in healthcare practice and usually should not be regarded as errors. Doctors may include information in the summary that was never documented, that was documented outside the considered context (in our case, only the BHC), or that was altered just prior to discharge. In this work, we considered the MIMIC data from a machine learning perspective and analyzed the effect of hallucinations in the training data on LLMs.
The labeling was carried out by two German medical students in their sixth year. They had completed their second state examination (USMLE Step 2 equivalent) and were working in the hospital. We utilized MedTator for annotation (He et al., 2022). For annotator training, we used twelve examples. Two examples were used to familiarize with the task and two times five examples were labeled separately and discussed for training. For the final labeling, the annotators worked independently and reached a consensus through discussion. To quantify the variability of the annotation, we determined manual and automatic statistics for agreement (see Appendix B).
4.4 Data-Centric Hallucination Reduction
We evaluated whether manually removing hallucinations from the training data can reduce hallucinations of LLMs. To this end, we tested Llama 70B fine-tuned on 100 examples and GPT-4 5-shot prompted with 5 random examples from the Original and the hallucination-free Cleaned data. We also included GPT-4 0-shot, which did not require any training data. Figure 11 contains examples for all five models. Due to limited manual annotation capabilities, no further models were included in this analysis. To determine the number of generated hallucinations, we chose 20 random contexts from MIMIC-IV-Note-Ext-DI-BHC-Anno and generated summaries with each of the five models. We annotated the resulting 100 summaries with our protocol yielding the hallucination dataset Hallucinations-Generated-DI. We also determined the amount of missing key facts and medical jargon using the annotation procedures of the qualitative evaluation (see Appendix F). The summaries for each context were permuted during the annotation process to prevent model identification.
@@ 4.4 Data-Centric Hallucination Reduction
Dataset / Number of Annotated Hallucinations per Type
@@
[…]
@@ 4.5 Quantitative Evaluation
We evaluated the performance of all models to compare them to existing work. We used the MIMIC-IV-Note-Ext-DI-BHC dataset, which contains 100,175 context-summary pairs. For training, we used 80,140 examples (80% of the data) for LED, 100 examples for Llama, and 5 examples for GPT-4 5-shot. We performed parameter and prompt tuning for all models (see Appendices C and D) and used 100 examples each for validation and testing from the remaining 20% of the held-out data.
To evaluate lexical overlap, we used the ROUGE F1 score (Lin, 2004). For similarity based on contextual embeddings, we reported BERTScore (Zhang et al., 2020a) for the default roberta-large (BERTScore) and microsoft/deberta-large-mnli (DeBERT) as recommended by the authors. [^4] See readme: https://github.com/Tiiiger/bert_score We did not utilize medical embeddings since the summaries should be written in layperson language. Lastly, we determined the SARI score for text simplification (Xu et al., 2016) and the number of generated words.
Figure 3: Patient summaries generated by Llama 70B fine-tuned on 100 Original and 100 Cleaned examples given the synthetic context in Figure 2 with annotated hallucinations according to our protocol. These are two of the five models included in the data-centric hallucination reduction experiments.
@@ 4.6 Qualitative Evaluation
We evaluated the quality of patient summaries generated by LLMs aligned with the highest quality data (Cleaned & Improved). Llama 70B was fine-tuned on 100 examples and GPT-4 5-shot was prompted with 5 random examples. We also included the original MIMIC summaries, LED-large fine-tuned on 20,942 MIMIC-IV-Note-Ext-DI-BHC-Anno examples, and GPT-4 0-shot (see examples in Figure 12). We generated summaries for the same 20 contexts from MIMIC-IV-Note-Ext-DI-BHC-Anno used for the hallucination-reduction experiments to reduce the annotation workload. Ratings were performed by the same two medical experts and the summaries were again permuted to prevent model identification. The annotators completed two training examples, and no agreement was sought as this was considered a subjective task. We combined both ratings for our analysis.
@@
We evaluated the summaries for Relevance, Consistency, Fluency, and Coherence on a 1-5 Likert scale (Fabbri et al., 2021) following previous work on medical summarization (Adams et al., 2022). We added the dimension Simplification. To obtain more reliable results, we defined subtasks to be performed before rating Relevance (subtask: labeling key facts in context and summary), Consistency (labeling hallucinations using our protocol), and Simplification (labeling medical jargon) and defined the meaning of each Likert rating based on these subtasks. Additional details on the subtasks and annotator instructions are given in Appendix F and Figures 8 and 9.
4.7 Automatic Hallucination Detection
We performed an evaluation of automated hallucination detection on two annotated datasets: 100 doctor-written summaries (Hallucinations-MIMIC-DI) and 100 LLM-generated summaries (Hallucinations-Generated-DI). This method could serve for the automatic cleaning of training data or post-hoc editing to improve faithfulness. We framed the task of automatic hallucination detection as a span classification problem. Given a context $C$ and a summary $S$, our objective was to identify all spans of hallucinations within the summary, denoted as $\mathcal{H}_{S}=\{H_{1},...,H_{n}\}$. Each hallucination span $H_{i}$ is defined by a tuple $(s,e,c)$, consisting of a start token, an end token, and a class (see Figure 2). We also explored class-agnostic hallucination detection without the class $c$.
We used a class-agnostic approach based on medical entities as a baseline (Cai et al., 2022; Adams et al., 2022). For entity recognition and linking, we utilized MedCat (Kraljevic et al., 2021) with UMLS semantic types from Adams et al. (2021). All entities that appeared in the summary but not in the context were considered hallucinations. We further enhanced this approach with SapBERT embeddings (Liu et al., 2021) to determine equivalent entities (Adams et al., 2023). The equivalence threshold for similarity was a hyperparameter for these methods and we determined its value based on ten additional examples labeled during annotator training. We tested GPT-4 for class-agnostic and class-aware hallucination detection. We designed a suitable prompt based on the aforementioned ten examples labeled during annotator training and utilized varying numbers of in-context examples and chain-of-thought (COT) prompting, detailed in Appendix E. Since our prompt design is based on the annotation protocol for general medical text, this approach might also prove useful for texts other than patient summaries. Evaluation was performed using partial matching with the nervaluate package (Segura-Bedmar et al., 2013).
[…]
MIMIC-IV-Note-Ext-DI-BHC (100,175 examples)
LED-large (80,140 ex.) $43.82$ $17.33$ $8.85$ $4.92$ $29.89$ $88.11$ $64.12$ $46.71$ $76.99$
Llama 2 7B (100 ex.) $38.36$ $12.66$ $5.13$ $2.24$ $24.73$ $85.68$ $60.23$ $44.12$ $73.13$
Llama 2 70B (100 ex.) $40.58$ $14.31$ $6.09$ $2.74$ $26.19$ $86.30$ $61.89$ $45.16$ $76.90$
GPT-4 5-shot (5 ex.) $38.80$ $10.78$ $3.55$ $1.12$ $21.98$ $86.67$ $61.30$ $42.88$ $131.86$
@@ 4.7 Automatic Hallucination Detection
Table 4: Quantitative evaluation of patient summary generation on MIMIC-IV-Note-Ext-DI-BHC dataset. We used the ROUGE F1-score for n-grams (R-n) and the longest common subsequence (R-L), BERTScore using roberta-large (BERT) and deberta-large-mnli (DeBERT), the SARI score, and the number of generated words. Additional performance results can be found in Table 11.
@@ 5.1 Hallucinations in MIMIC and Generated Patient Summaries
Two medical experts labeled 100 doctor-written patient summaries (Hallucinations-MIMIC-DI) and found 286 hallucinations (see Table 3). The most prevalent label was the generic word unsupported (n=76) followed by condition unsupported (n=52), and time unsupported (n=35). Hence, there is a significant amount of unsupported data in MIMIC when using the BHC as context. For the 100 generated patient summaries (Hallucinations-Generated-DI), they found a total of 114 hallucinations and word unsupported was the most common label (see Table 3). Agreement statistics show that annotators agreed on 1.55 of 2.86 annotations for MIMIC and 0.67 of 1.14 annotations for the generated summaries (see Table 8) emphasizing some variability during labeling. The annotation of MIMIC summaries took 30 hours for each expert and 6 hours for the agreement. For the generated summaries, annotation took less time, with 20 hours for labeling and 4 hours for the agreement, since there were only 20 different patient contexts.
@@ 5.2 Data-Centric Hallucination Reduction
The results for GPT-4 5-shot prompted with Original data were significantly better than for Llama 70B fine-tuned on Original with 0.70 versus 2.60 hallucinations and 0.93 versus 3.77 missing key facts per summary (see Table 2). Data-centric hallucination reduction (training on Cleaned), showed a substantial effect for Llama 70B, reducing hallucinations from 2.60 to 1.55 per summary while maintaining the same number of key facts. Figure 3 displays generations of both models with hallucination annotations. Llama 70B trained on Original introduced unsupported adjectives “mild” and “other”, as well as an unsupported metformin dosage. Both models incorrectly recommended continuing antibiotics at home. Apart from “mild”, all hallucinations can be considered reasonable advice and are commonly found in patient summaries. For GPT-4 5-shot, the difference of 0.70 (Original) to 0.40 (Cleaned) hallucinations is only marginal, while retaining key facts. The use of medical jargon and the length of the generations did not change substantially. GPT-4 0-shot exhibited a few hallucinations (0.45) and the lowest number of missing key facts (0.82) and medical jargon (0.70), although the generations were longer than for 5-shot. Table 3 shows the types of hallucination for each model. While Llama 70B occasionally generated unsupported conditions and locations, this was rare for GPT-4. The hallucinations produced by GPT-4 were mostly labeled as unsupported word.
Figure 4: Qualitative evaluation of Llama 70B fine-tuned on all 100 examples of Cleaned & Improved and GPT-4 5-shot prompted with 5 random examples of Cleaned & Improved. We compared them to the original MIMIC summaries, LED-large fine-tuned on MIMIC-IV-Note-Ext-DI-BHC-Anno, and GPT-4 0-shot. Two medical experts evaluated 20 summaries from each of the five models.
@@ 5.3 Quantitative Evaluation
Table 4 contains the performance evaluation on the MIMIC-IV-Note-Ext-DI-BHC dataset. Overall, the LED model performed the best, followed by Llama and GPT-4. Only for BERT and DeBERT, Llama and GPT-4 achieved similar performance. This behavior is consistent with the amount of training data, which is much less for Llama and GPT-4 5-shot. Notably, GPT-4 0-shot achieved considerable performance even without using any training data.
@@ 5.4 Qualitative Evaluation
The qualitative results are visualized in Figure 4. The Consistency results for doctor-written (MIMIC) summaries should be interpreted with caution because these summaries were artificially considered only with the restricted context of the BHC. The original MIMIC examples, LED-large trained on all data, and Llama 70B fine-tuned on 100 Cleaned & Improved examples showed similar results. For Relevance, around half of the examples received a rating of two or higher, indicating that at most three key points were missing. The Simplification, Fluency, and Coherence all received a median rating of four for all three approaches. Llama 70B performed better for Consistency with a median rating of four compared to three for the original MIMIC and LED-large summaries. GPT-4 performed significantly better across all five dimensions. The difference in Consistency compared to Llama 70B is in line with the lower hallucination rate of GPT-4 observed earlier (see Table 2). Also, GPT-4 achieved high Relevance with a median of four, indicating one missing key fact per summary, and a median of five for simplification, fluency, and coherence. There was only a slight difference between GPT-4 5-shot and GPT-4 0-shot. GPT-4 5-shot performed better for Consistency, i.e., including all key facts, while GPT-4 0-shot performed better for Relevance and Simplification.
Model H.-MIMIC-DI H.-Generated-DI
@@
Table 5: Results of automatic hallucination detection on 100 doctor-written (H.-MIMIC-DI) and 100 LLM-generated (H.-Generated-DI) patient summaries. GPT-4 clearly outperformed the baselines using medical UMLS entities recognized by MedCat.